# A Bayesian Perspective on Evidence for Evolving Dark Energy

### v1

@@ II Methods
We analyse DESI DR2 BAO data [desi2025, BAOData] combined with Planck 2018 CMB (CamSpec likelihood [CamSpec2021]) and Type Ia supernovae (Pantheon+ [Scolnic2018], Union3 [rubin2025unionunitycosmology2000], DES-Y5 [descollaboration2025darkenergysurveycosmology]). We use the PolyChord [Handley2015PolychordI, Handley2015PolychordII] nested sampling algorithm via Cobaya [Torrado2021Cobaya, cobayaascl] and CAMB [Lewis:1999bs]. For CDM, we adopt DESI’s priors (, with ), except setting a 0.06 eV lower bound on neutrino mass from oscillation experiments, versus DESI’s 0 eV lower limit.
@@ III Results
Table 1 compares the DESI collaboration’s frequentist tests with our Bayesian model comparison for CDM versus CDM. From a Bayesian perspective, the DESI DR2 data alone penalises the complexity of the CDM model, favouring the simpler CDM with a log-Bayes factor of . This preference for CDM remains when CMB data are added. The combination of DESI DR2 + CMB, which the DESI collaboration highlights as a key result, yields in our analysis. This stands in contrast to the frequentist finding of a 3.1 preference for CDM, setting the stage for a methodological comparison. A similar pattern is observed for DESI DR1 data, which also favours CDM both individually and in combination with the CMB.
@@ IV Conclusions
Using established cosmological data analysis methodology, we provide a principled benchmark for the Bayesian evidence of CDM across multiple probes. This exercise reveals a discrepancy between our Bayesian results and the published frequentist test results, of a magnitude that cannot reasonably be ascribed to philosophical distinctions between the paradigms. We suggest that reliance on asymptotic formulae alone may be a contributing factor, given the modest effective dimensionality of the DESI BAO data set. Although both approaches penalize the additional flexibility of CDM, they do so via distinct mechanisms. Importantly, our finding that DESI BAO + CMB data are statistically consistent with CDM is independently corroborated by the geometric analysis of Efstathiou [Efstathiou2025BAO], which offers an alternative, principled framework for assessing model consistency. The convergence of Bayesian model comparison, targeted tension diagnostics, and independent geometric analyses provides a robust statistical basis for the conclusion that claims of evidence for evolving dark energy are overstated and largely driven by tension with a single external data set. All analysis products are available via the unimpeded framework [UnimpededPaper, UnimpededSoftware].

### v2

@@ II Methods
We analyse DESI DR2 BAO data [desi2025, BAOData] combined with Planck 2018 CMB (CamSpec likelihood [CamSpec2021]) and Type Ia supernovae (Pantheon+ [Scolnic2018], Union3 [rubin2025unionunitycosmology2000], DES-Y5 [descollaboration2025darkenergysurveycosmology]). We use the PolyChord [Handley2015PolychordI, Handley2015PolychordII] nested sampling algorithm via Cobaya [Torrado2021Cobaya, cobayaascl] and CAMB [Lewis:1999bs]. For CDM, we adopt DESI’s priors (, ).
@@ III Results
Table 1 compares the DESI Collaboration’s frequentist hypothesis tests with our Bayesian model comparison for CDM versus CDM. From a Bayesian perspective, the DESI DR2 data alone penalises the complexity of the CDM model, favouring the simpler CDM with a log-Bayes factor of . This preference for CDM remains when CMB data are added. The combination of DESI DR2 + CMB, which the DESI Collaboration highlights as a key result, yields in our analysis. This stands in contrast to the frequentist finding of a 3.1 preference for CDM, setting the stage for a methodological comparison. A similar pattern is observed for DESI DR1 data, which also favours CDM both individually and in combination with the CMB.
@@ IV Conclusions
Using established cosmological data analysis methodology, we provide a principled benchmark for the Bayesian evidence of CDM across multiple probes. This exercise reveals a discrepancy between our Bayesian results and the published frequentist hypothesis test results, of a magnitude that cannot reasonably be ascribed to philosophical distinctions between the paradigms. We suggest that reliance on asymptotic formulae alone may be a contributing factor, given the modest effective dimensionality of the DESI BAO data set. Although both approaches penalize the additional flexibility of CDM, they do so via distinct mechanisms. Importantly, our finding that DESI BAO + CMB data are statistically consistent with CDM is independently corroborated by the geometric analysis of Efstathiou [Efstathiou2025BAO], which offers an alternative, principled framework for assessing model consistency. The convergence of Bayesian model comparison, targeted tension diagnostics, and independent geometric analyses provides a robust statistical basis for the conclusion that claims of evidence for evolving dark energy are overstated and largely driven by tension with a single external data set. All analysis products are available via the unimpeded framework [UnimpededPaper, UnimpededSoftware].